Kevin earned $165 interest for 1 year on an investment of
$1500. At the same rate, what amount of interest would be earned by an investment of $2500?

Answer by checkley71(8403)

165/1,500=x/2,500 cross multiply.
1,500x=165*2,500
1,500x=412,500
x=412,500/1,500
x=275 answer.